OpenStudy (anonymous):

find y'. y=ln e^x/e^x-1

OpenStudy (anonymous):

is this written correctly: \[y = \ln \frac{ e^{x} }{ e^{x} - 1 }\]

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

\[\ln u = \frac{ du }{ u }\] this is a simplified version of it use the chain rule here where u = e^x / (e^x - 1) and du = the derivative of that

OpenStudy (anonymous):

I tried that but got stuck trying to work out the derivative of e^x and e^x-1

OpenStudy (anonymous):

\[\frac{ d }{ dx } \frac{ e^{x} }{ e^{x} - 1 }\] i was taught in high school "down d up - up d down all over down squared" just means "denominator * derivative of numerator - numerator * derivative of denominator, all divided by the square of the denominator" in this case, the derivative of the numerator is e^x, and the derivative of the denominator is also e^x so we have \[\frac{ d }{ dx }\frac{ e^{x} }{ e^{x} - 1 } = \frac{ (e^{x} - 1) * e^{x} - e^{x} * e^{x} }{ (e^{x} - 1)^{2} }\]

OpenStudy (anonymous):

that simplifies to:

OpenStudy (anonymous):

\[-\frac{ e^{x} }{ (e^{x} - 1)^{2} }\]

OpenStudy (anonymous):

do you mind breaking down how you go t that answer. please :)

OpenStudy (anonymous):

that's just the derivative of e^x / (e^x - 1) you use the quotient rule for derivatives

OpenStudy (anonymous):

and we said u = e^x / (e^x - 1), and du, we just solved for, du = -e^x / (e^x - 1)^2 therefore the derivative of ln u = du/u = \[\frac{ -e^{x} (e^{x} - 1) }{ e^{x}(e^{x} - 1)^{2} } = \frac{ -1 }{ e^{x} - 1 } = \frac{ 1 }{ 1 - e^{x} }\]

OpenStudy (anonymous):

okay i think i got it . thank you so much